# SynLlama: Generating Synthesizable Molecules and Their Analogs with Large Language Models

### v1

@@ 4 Discussion
Motivated by recent advances in LLMs for chemistry54, 52, 38, 53, we aimed to leverage data-efficient supervised fine-tuning (SFT) techniques to transform the general-purpose Meta Llama 3 into SynLlama, an LLM-based generator capable of sampling synthesizable molecules or close analogs and deducing their synthetic routes given desired target molecules. Throughout the study, we successfully show that SynLlama can effectively explore a custom-defined chemical search space composed of around 230,000 Enamine building blocks (BBs) and 91 well-validated organic reactions (RXNs), after it has been fine-tuned on synthetic pathway data sampled from this specified chemical space. What’s more, despite utilizing nearly two orders of magnitude fewer synthetic pathways, SynLlama exhibits strong performance in key drug discovery tasks compared to existing models. Specifically, we have demonstrated that SynLlama can effectively aid in various stages of drug discovery that include synthesis planning, synthesizable analog generation for de novo molecules, and local hit expansion.
@@
Because SynLlama is built off of a LLM, as opposed to building it as a chemical language model from scratch, it offers a number of unique advantages and possibilities for further improvement. For example when generating our fine-tuning data, we sampled from the predefined chemical search space of 230K Enamine building blocks (BBs) and 91 reaction templates (RXNs), but we didn’t embed these extensive requirements in the context window of the LLM. As a result for our largest dataset with 2 million synthetic pathways, each BB was seen by the model only a few dozen times, while each RXN template appeared hundreds of thousands of times. Consequently while SynLlama efficiently memorizes the 91 allowed RXNs, it only captures the distribution of Enamine BBs such that SynLlama is able to extrapolate to unseen yet purchasable building blocks. This generative ability surpasses other existing methods, which typically formulate the synthesis pathway generation as a Markov decision process within a defined chemical space.
@@
In addition to its ability to extrapolate outside the training chemical space, the underlying Llama-3.2-1B used by SynLlama is relatively small and more predictive power would be expected if we train on larger LLMs with more data and compute power. However, we observe that a smaller LLM with fewer parameters can be turned into an expert model for complex tasks after SFT with sufficient data. This opens up opportunities to employ smaller expert models across various chemical tasks, benefiting from faster inference speeds, which can make these models more desirable. Moreover, optimal hyperparameters such as temperature and top-p can differ between training, for which generating raw outputs with good chemistry comprehension is important, and inference for enhancing sampling capability as a generative model to explore chemical space more effectively, thereby improving diversity of the generated analogs. This property is especially desirable in tasks that require extensive exploration, such as the hit expansion example we demonstrate. Another exciting success that we observe is SynLlama’s utility when it’s coupled with another generative model. The fact that we see a retainment in docking score and distributional shift in SA score suggests that SynLlama can effectively serve as a post-processor for other de novo generative models to ensure more synthesizable compounds that have explicit reaction pathways.
@@
Among the numerous opportunities for LLMs in areas of chemistry such as drug discovery, their natural language capabilities and recent advances in their reasoning abilities can enable users without coding expertise to interact with LLM-based tools directly, effectively bridging the gap between computational methods and experimental research. We imagine that expert users (as opposed to developers) can employ additional prompt engineering and SFT data that would allow for incorporation of more realistic factors than that explored here, such as synthesis cost, reaction conditions, improved selectivity, need for protection factors at certain reaction steps, etc. into more detailed and powerful synthesis planning.
@@ 5 DATA AND CODE AVAILABILITY
All the codes for SynLlama and SFT dataset creation are provided in a public accessible GitHub repository: https://github.com/THGLab/SynLlama under MIT License.

### v2

@@ 4 Discussion
Motivated by recent advances in LLMs for chemistry54, 52, 38, 53, we aim to leverage data-efficient supervised fine-tuning (SFT) to transform the general-purpose Meta Llama 3 into SynLlama, an LLM-based generator capable of proposing synthesizable molecules and deducing synthetic routes for target molecules or their close analogs. Throughout the study, we successfully show that SynLlama can effectively explore a custom-defined chemical search space composed of around 230,000 Enamine building blocks (BBs) and well-validated organic reactions (RXNs), after it has been fine-tuned on synthetic pathway data sampled from this specified chemical space. What’s more, despite utilizing nearly two orders of magnitude fewer synthetic pathways, SynLlama exhibits strong performance in key drug discovery tasks compared to existing models. Specifically, we have demonstrated that SynLlama can effectively aid in various stages of drug discovery that include synthesis planning, synthesizable analog generation for de novo molecules, and local hit expansion.
@@
Because SynLlama is built on a general-purpose LLM instead of training from scratch48, 49, 50, it offers a number of unique advantages and possibilities for further improvement. For example, when generating our fine-tuning data, we sampled from the predefined chemical search space of 230K Enamine building blocks (BBs) and two sets of reaction templates (RXNs), but we did not embed these extensive requirements in the context window of the LLM. As a result, for our largest dataset with 2 million synthetic pathways, the model only saw each BB a few dozen times, while each RXN template appeared hundreds of thousands of times. Consequently, while SynLlama efficiently memorizes the allowed RXNs, it only captures the distribution of Enamine BBs, which enables SynLlama to extrapolate to unseen yet purchasable building blocks outside of Enamine. This generative ability surpasses other existing methods, which typically formulate the synthetic pathway generation as a Markov decision process within a defined chemical space.
@@
In addition to its ability to extrapolate outside the training chemical space, the underlying Llama-3.2-1B used by SynLlama is relatively small and more predictive power would be expected if we train on larger LLMs with more data and compute power. However, we observe that a smaller LLM with fewer parameters can be turned into an expert model for complex tasks after SFT with sufficient data. This opens up opportunities to employ smaller expert models for various chemical tasks, benefiting from faster inference speeds, which can make these models more desirable. Moreover, optimal hyperparameters like temperature and top-p can vary between training and inference phases, depending on the downstream tasks. During inference, most valid raw outputs are generated under relatively low temperature and top-p settings. However, when the model is paired with reconstruction algorithms that require less strict adherence to reaction chemistry, higher temperature and top-p values can be used. This allows for a broader exploration of the Enamine chemical space, enabling the generation of more diverse and relevant analogs. This property is especially desirable in tasks that require extensive exploration, such as the hit expansion example we demonstrate. Another exciting success is that when SynLlama is paired with another generative model, the generated analogs can maintain docking scores while exhibiting a shift in synthetic accessibility scores. This result suggests that SynLlama can serve effectively as a post-processor for other de novo generative models, ensuring the production of more synthesizable compounds with clear reaction pathways.
@@
Among the numerous opportunities that LLMs bring to the field of drug discovery, their natural language capabilities and recent advancements in reasoning are the most exciting features that allow users without coding expertise to interact directly with the models, effectively bridging the gap between computational methods and experimental research. We envision that expert users can employ prompt engineering and fine-tuning data to incorporate more realistic factors than those explored here. For instance, medicinal chemists could fine-tune LLMs within this generalizable SFT framework with building blocks and reaction templates of their own choice. In addition, they can consider synthesis cost, reaction conditions, improved selectivity, and protection factors at specific reaction steps for more detailed and powerful synthesis planning. We see our work as an initial attempt to demonstrate the effectiveness of LLMs in real experimental research, encouraging further studies for better utilization of these models.
@@ 5 DATA AND CODE AVAILABILITY
All the codes and data for SynLlama workflow are provided in a public accessible GitHub repository: https://github.com/THGLab/SynLlama under MIT License.